OpenStudy (anonymous):

How to solve cot(-45*)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

how to solve?

OpenStudy (anonymous):

cot is just cos/sin, right? Now you have cos(-45) --------- sin(-45) Now the cool thing about cos, is that it is an even function. Which means we have now: 1/sqrt(2) ---------- -(1/sqrt(2) Which simplifes to: 1 ---- -1 which equals: -1

OpenStudy (anonymous):

sorry for not checking the notifications ;P. but just one question there, um i get the part where you did cos(-45)/sin(-45). how did 1/root(2) change to 1 then -1?

OpenStudy (anonymous):

\[\cos(-45)=\frac{1}{\sqrt{2}}\] \[\sin(-54)=-\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

typo

OpenStudy (anonymous):

\[\sin(-45)=-\frac{1}{\sqrt{2}}\]

OpenStudy (anonymous):

take cosine over sine, get -1

OpenStudy (anonymous):

okay okay now that explained it! THANKS ! XD!!

OpenStudy (anonymous):

\[\frac{\frac{1}{\sqrt{2}}}{-\frac{1}{\sqrt{2}}}=-1\]